# Prompt Selection Matters: Enhancing Text Annotations for Social Sciences with Large Language Models

### v1

@@ Abstract
Large Language Models have recently been applied to text annotation tasks from social sciences, equalling or surpassing the performance of human workers at a fraction of the cost.
However, no inquiry has yet been made on the impact of prompt selection on labelling accuracy.
@@ 1 Introduction
Throughout the social sciences, many research questions are answered through annotation and classification of large volumes of text, such as tweets or Facebook comments. Researchers may be interested in knowing, for instance, how politically slanted (liberal vs. conservative) a claim or headline is; how emotional or hostile its tone is, or which one of the basic emotions it reflects (sadness, joy, anger, etc) [BLHJ16, GL12, SM10, RBOP24]. Text annotation so far had to be performed either by human experts or by unskilled crowd workers, depending on the nature of the task. As a result, it was typically costly and time-consuming; crowd workers are also likely to mislabel the data.
The recent progress of Large Language Models (LLMs) has opened new avenues and could revolutionize text mining by allowing huge volumes of text data to be analyzed in an unsupervised way in a matter of minutes [HCvH24, GAK23, Tör23, WR23]. Early results suggest that LLMs can perform extremely well, extremely fast and at almost no financial cost, with levels of accuracy rivalling those of experts and superior to those of unskilled workers. Moreover, these results are achieved by off-the-shelf, general purpose models, that do not need any specialized training, unlike some pre-existing machine learning-based techniques.
Nonetheless, crucial aspects of the automatic annotation of text data using LLMs have not been studied yet. In particular, earlier studies did not consider the importance of prompt choice: they used simple hand-crafted prompts, such as “Does the following message express liberal or conservative views?” or “Is this tweet pro-life or pro-choice?”, to annotate their corpus. However, it has been observed outside the context of text annotation that applying distinct yet similar versions of a given prompt to certain tasks can result in large differences in accuracy, of the order of more than [KGR+22, BMR+20]. Going from to of mislabelled data can greatly impact the quality of a study’s conclusions, especially if the classification errors are biased (e.g., if all mislabelled tweets express conservative views).
Our main objectives in this paper are threefold. First, to help social scientists understand and adopt automatic text annotation using LLMs by providing a clear illustration of it. Second, to investigate the importance of prompt selection on performance, and to raise awareness about the issue of performance variability. Third, to explain to social scientists how to apply state-of-the-art prompt optimization methods used in the wider LLM community [ZMH+22, YWL+24] to their own classification tasks.
@@
We describe the principle of automatic prompt optimization, and provide a simple implementation of the method.
We investigate and quantify the impact of prompt selection on accuracy across a range of standard classification tasks in social sciences using LLMs. To that end, we compare both hand-crafted prompts and automatically optimized prompts.
We conclude that apparently similar prompts yield greatly varied accuracy levels. We also observe that automatic prompt optimization yields consistently good performance and beats prompt-crafting heuristics (such as using Chain of Thoughts prompts, see below) on most tasks.
Finally, we provide the community with a simple and efficient way to label their datasets using LLMs coupled with either their own hand-crafted prompts or with a prompt optimization algorithm: our browser-based service https://prompt-ultra.github.io/.
All of the code used will be made available on GitHub.
@@ 2 Automatic text labelling using LLMs
State-of-the-art general purpose language models, such as GPT-4 [Ope23], have now achieved human-like performance on a variety of tasks. In particular, it has been recently demonstrated that they can be applied off-the-shelf, i.e., without needing any specialized training, to a variety of text-labelling tasks needed in social sciences. Among other examples, they have been shown to reach around of accuracy111It is useful to note in passing that while it is desirable to maximise accuracy at classification tasks, it is generally impossible to reach 100% in virtue of the fact that some statements are intrinsically ambiguous and thus impossible to classify even by professional raters. on negativity detection in tweets and news articles [HCvH24], around of accuracy on stance detection in tweets (largely outperforming untrained crowd workers) [GAK23], and more than of accuracy on political orientation in tweets [Tör23].
@@
Typing each prompt by hand in the browser interface would be tedious. Instead, researchers would typically send such prompts through the API using e.g., a simple Python script222Such a script can be found in our code base., allowing thousands of statements to be classified in a matter of minutes. Each message would have to be processed in a fresh instance of the client chat, to avoid a progressive biasing of the model by the conversation history. A small percentage of answers might not be correctly formatted (the LLM might respond “The message is conservative” or “Conservative.” or “right-wing” instead of “conservative”), in which case one would simply have to ask the chatbot again. To get an estimate of the accuracy of the labels thus produced, the researcher should ideally have a small subset of the dataset be annotated by experts (typically a few hundreds to a few thousand items of text), so that the labels provided by the LLM can be compared to those provided by experts, which serve as ground truth for the task. The degree of agreement between the two sets of labels can then serve as an estimate of the accuracy of the LLM’s annotations.
The perks of using LLMs to classify text, compared to manual annotation by crowd workers or experts, are immense. It is easy to implement, fast and inexpensive (thousands of short or medium length texts can be labelled in a few minutes for less than 10$). It also outperforms earlier machine learning-based methods without requiring the training of a specialized model, which can be time-consuming and requires technical skills.
@@ 2.1 Limitations of automatic text labelling
As for any method, automatic text annotation using LLMs suffers from some limitations, some of which have not been extensively discussed in the literature. Though early studies are very encouraging, the range of tasks that have been studied so far remains rather limited; the vast majority of them consists in attributing one of a handful of labels to short- to medium-length texts (most of which being social media posts or news headlines). Moreover, some preliminary results find that accuracy tends to be lower in languages other than English [HCvH24], which is coherent with general observations regarding LLMs’ performances. LLMs are also known to suffer from certain biases, including political ones [LJW+22, MPNR24, ZLS+24]; those could result in systematic biases on certain tasks that would have a greater impact on the research performed on the annotated dataset than the raw accuracy might suggest. As an example, consider an LLM tasked with labelling a set of tweets as Republican-leaning or Democrat-leaning ; if all mislabelled tweets are Republican tweets erroneously labeled as Democrat, the conclusions drawn from the dataset might be wrong, despite the overall accuracy being high.
@@
The issue can also be exacerbated by conscious moderating efforts by the LLM’s creators: various techniques are in place to ensure that the answers provided by LLM-based chat agents avoid using stereotypes susceptible to be seen as offensive, which could negatively affect their performance on some politically sensitive tasks, such as race.
@@
Other problems arise not from the intrinsic nature of LLMs, but rather from the way in which they are currently being trained and deployed. State-of-the-art LLMs, such as GPT-4, are trained on immense datasets of text scrapped from the web, and regularly updated in a similar fashion. Beyond the possible privacy issues, this has crucial implications regarding the performance of LLMs: they will often perform better on data on which they have been trained than on yet unseen data (e.g., data that was produced after their training). As a result, the same LLM might be much better at classifying a certain set of tweets than a similar yet more recent set which it has not yet “seen”. As the exact nature of the training data is typically not shared with the general public, such phenomena are very hard to predict - see some of our observations in Section 4. In particular, we suspect that some of the most impressive accuracy scores achieved by LLMs on annotation tasks and reported in the literature might be partially explained by this. This problem, in conjunction with the regular updating of models, creates important issues of replicability: results might vary greatly between apparently similar tasks, or when applying different versions of the same model to the same task. This can be partially offset by using “frozen” LLMs, i.e. by keeping a copy of a certain version of an LLM at a certain point in time and using it on all tasks.
@@ 3 Experiments
We evaluate various hand-crafted prompts and automatically optimized prompts (see Subsections 3.2 and 3.3 below) on a range of datasets and classical social sciences tasks using OpenAI’s GPT-3.5 Turbo’s API [YCX+23] (see our code base for details). Our objective is twofold: first, we want to measure the impact of the precise formulation of the prompt on GPT’s accuracy by comparing performance across various prompts. The goal here is not to find some optimal prompt crafting technique that would systematically outperform all others (such panacea is unlikely to exist, as suggested by our experimental results further below); in particular, we do not claim to have examined every possible prompt-crafting trick. Rather, we only want to see whether two reasonable, semantically similar prompts can result in significantly different accuracies on a given task. Second, we want to check whether automatic prompt optimization can help achieve consistently good (though not necessarily optimal) results on all tasks without the need for manual tweaking by the experimenter.
@@ 3.1 Datasets and tasks
A dataset of news articles from major US newspapers [BDSMGN20], with labels representing their political inclination (left, center333Note that the centrist label describes articles that are biased towards a centrist political ideology, and not articles that lack political bias. or right), from which we randomly sample articles.
@@ 3.2 Hand-crafted prompts
It has been observed outside the context of automatic text annotation that various tricks can help increase LLMs’ performance. As an example, simply adding “Let’s think step by step” at the end of a prompt can result in impressive improvements [KGR+22]. Though there is no systematic theory of “good prompting”, many such tricks have been catalogued (see e.g., [pro]). We have applied several of these standard techniques to craft five different prompts for each task; below are examples from the TweetEval - hate task. The other prompts can be found in the accompanying code.
@@ 3.3 Automatic prompt optimization (APO)
The method is both conceptually simple, and easy to implement; its main a priori downside compared to using hand-crafted prompts is the increased number of calls to the chatbot API required, though the costs remain almost negligible. It can be tested on our free browser-based service, which we further describe in Appendix A.
@@ 4 Results and discussion
| Accuracies | |||||||
| TE-hate | TE-emotion | TE-sent | TE-off | TML-sent | AS-pol | LibCon | |
| Simple | 57.0 | 60.5 | 71.4 | ||||
| Explanations | |||||||
| Examples | 76.9 | ||||||
| Roleplay | 65.5 | 47.1 | 70.7 | ||||
| CoT | |||||||
| APO | |||||||
@@
The accuracy444We also recorded F1 scores, but do not report them, as they are very correlated with the accuracy and hence somewhat redundant. on each task of our five types of prompts and of the best prompt generated using automatic prompt generation are reported in Table 1 We observe that the choice of prompt has a significant impact on performance for most tasks: depending on the task, the worst prompt results in between (LibCon) and (TE-off) more errors than the best prompt. This confirms our claim that careful prompt selection is crucial.
We also note that there is no "miracle" hand-crafted prompt that consistently outperforms all the others. Some heuristics and specific wordings work well on certain tasks, and poorly on others. This is in line with observations made in the literature; as an example, adding “Let’s think step by step” at the end of a prompt is shown to outperform adding the semantically similar “Let’s work this out in a step by step way to be sure we have the right answer” for the task studied in [YWL+24] (see Table 1), while the converse is true in [ZMH+22] (Table 7).
@@
We now turn to the specific performances of Automatic prompt optimization (APO). APO beats hand-crafted prompts at all tasks but two; among those two, it is second best for TE-sent (with a score nearly equal to that of the best hand-crafted prompt), and is roughly at the level of the median of the hand-crafted prompts for TE-off555Note that TE-off has the smallest test set of all the tasks, and as a result the measured accuracies suffer from a greater variance.. This shows that automatically optimizing prompts results in consistently good performance without the need for prior testing of various hand-crafted prompts on a subset of the dataset of interest.
Interestingly, and in line with the remarks of the previous paragraph, the final prompts generated using APO do not differ much from the ones used to start the optimization process, despite the often large gap between their associated accuracies666We report the best prompt returned by the optimization process for each task in the Appendix.. For TE-hate, for example, the initial prompt is “Classify the following message as hateful if it contains hate speech or non-hateful if it does not contain hate speech. Output only “hateful” or “non-hateful” without quotes” and the automatically optimized prompt is “Check for hate speech in the following message to determine if it is hateful, then classify it as either "hateful" or "non-hateful"”.
@@
| Train set vs test set | ||||||
| Simple | Explanations | Examples | Roleplay | CoT | APO | |
| TE-hate, train | ||||||
| TE-hate, test | ||||||
| TE-emotion, train | ||||||
| TE-emotion, test | ||||||
@@
We have observed a curious phenomenon: as shown in table 2, all prompts result in much higher accuracy when tested on the training set of the TE-hate dataset than on its test set (where “training” and “test” refer to the split made by the creators of the dataset [BBF+19]). This is a priori surprising, as both subsets should be drawn from the same distribution, and as such our prompts should not perform any better on the training set than on the test set777Note that there is no issue of overfitting here, as the prompts were not “trained” in any way on what we call the training set in this context.. Our explanation is that the training set of TE-hate is part of the enormous amount of data on which GPT-3.5 Turbo, the model which we used, has been trained; as a result, it performs better on it than on the test set (which was either not included, or included in a different fashion). As the exact data on which the model was trained was not made public, we cannot confirm this hypothesis with absolute certainty. As a result, we can say that ChatGPT is “cheating” when labelling the training set of TE-hate, in the sense that its performance is superior to what it would be on a similar dataset to which it did not have access during its training (e.g. data that was produced after its training). This explains why we used the test splits of our datasets whenever possible, in the hope that as (we assume it to be the case) for TE-hate, they have not been included in the training set of the model, making for a fairer assessment of the capabilities of the method.
@@
This problem makes it hard to predict future performances based on past experiments, as we have no guarantee as to what was and was not used in the models’ training. Note for example that the same gap in performance between the training set and the test set is not observable for TE-emotion. Ideally, one should only test the models on data that was produced after their training. This could however prove difficult and overly restrictive: curating useful datasets is a challenging and time-consuming task in itself, and most datasets available online are at least a few years old, thereby predating all last generation LLMs.
@@
As a side note, this phenomenon could explain why we have observed slightly lower accuracies on average than those reported in recent literature [HCvH24, GAK23, Tör23] on similar tasks: we suspect that the authors of these articles might not have been aware of this pitfall, and have taken no measures to circumvent it. It could also be due to intrinsic differences in difficulty between the datasets used, as we tried to include both easy and hard datasets for diversity (in particular, it has been noted before, e.g. in [HCvH24], that performance drops for very long texts, such as the articles from the AS-pol dataset). We cannot verify this hypothesis, as (most of) the datasets used in the references that we cite are not available anymore due to a policy change from Twitter.

### v2

@@ Abstract
Large Language Models have recently been applied to text annotation tasks from social sciences, equating or surpassing the performance of human workers at a fraction of the cost.
However, no inquiry has yet been made of the impact of prompt selection on labelling accuracy.
@@ 1 Introduction
Throughout the social sciences, many research questions are answered using annotation and classification of large volumes of text, such as tweets or Facebook comments. Researchers may be interested in knowing, for instance, how politically slanted (liberal vs. conservative) a claim or headline is; how emotional or hostile its tone is, or which one of the basic emotions it reflects (sadness, joy, anger, etc) [BLHJ16, GL12, SM10, RBOP24]. Text annotation so far had to be performed by humans–experts, research assistants, or unskilled crowd workers who are likely to mislabel the data [Li24]–or using machine learning models trained via supervised learning specifically for the task at hand. As a result, it was typically costly and time-consuming.
The recent progress of Large Language Models (LLMs) has opened new avenues and could revolutionize text mining by allowing huge volumes of text data to be analyzed in an unsupervised way in a matter of minutes [HCvH24, GAK23, Tör23, WR23, DKG22]. Early results suggest that LLMs can perform extremely well, extremely fast and at a reasonable cost, with levels of accuracy rivaling those of experts and superior to those of unskilled workers. Moreover, these results are achieved by off-the-shelf, general purpose models, that do not need any specialized training, unlike some pre-existing machine learning-based techniques.
Nonetheless, crucial aspects of the automatic annotation of text data using LLMs have not been studied yet. In particular, few studies have considered the importance of prompt choice in this context: simple hand-crafted prompts, such as “Does the following message express liberal or conservative views?” or “Is this tweet pro-life or pro-choice?”, are frequently used to annotate corpora. However, it has been observed outside the context of text annotation that applying distinct yet similar versions of a given prompt to certain tasks can result in large differences in accuracy111Similar differences can be observed when using other performance metrics, such as F-scores., of the order of more than [KGR+22, BMR+20]; similar findings within the context of text annotation are reported in [WR23]. Going from to of mislabelled data can greatly impact the quality of a study’s conclusions, especially if the classification errors are biased (e.g., if all mislabelled tweets express conservative views).
Our main objectives in this paper are threefold. First, to help social better social scientists understand automatic text annotation using LLMs by providing a clear illustration of how it works. Second, to demonstrate the importance of prompt selection on performance, and to raise awareness about the issue of performance variability. Third, to explain to social scientists how to apply state-of-the-art prompt optimization methods used in the wider LLM community [ZMH+22, YWL+24] to their own classification tasks.
@@
We describe the principle of automatic prompt optimization, as well as a simple implementation of the method.
To the social science community, we provide a website that enables researchers to efficiently label their own datasets using LLMs using either their own hand-crafted prompts or a prompt optimization algorithm. This browser-based service can be found at https://prompt-ultra.github.io/.
We investigate and quantify the impact of prompt selection on performance across a range of standard classification tasks in social sciences using LLMs. To that end, we compare both hand-crafted prompts and automatically optimized prompts.
We conclude that apparently similar prompts in natural language yield greatly varied performance levels. We also observe that automatic prompt optimization yields consistently good performance and beats prompt-crafting heuristics (such as using Chain of Thoughts prompts, see below) on most tasks.
All of the code used to perform the experiments will be made available on GitHub.
@@ 2 Automatic text labelling using LLMs
State-of-the-art general purpose language models, such as GPT-4 [Ope23], have now achieved human-like performance on a variety of tasks. In particular, it has been recently demonstrated that they can be applied off-the-shelf, i.e., without needing any specialized training, to a variety of text-labelling tasks needed in social sciences. Among other examples, they have been shown to reach around of accuracy222It is useful to note in passing that while it is desirable to maximize accuracy at classification tasks, it is generally impossible to reach 100% in virtue of the fact that some statements are intrinsically ambiguous and thus impossible to classify even by professional raters. on negativity detection in tweets and news articles [HCvH24], around of accuracy on stance detection in tweets (largely outperforming untrained crowd workers) [GAK23], and more than of accuracy on political orientation in tweets [Tör23].
@@
Typing each prompt by hand in the browser interface would be tedious. Instead, researchers would typically send such prompts through the API using e.g., a simple Python script333Such a script can be found in our code base., allowing thousands of statements to be classified in a matter of minutes. Each message would have to be processed in a fresh instance of the client chat, to avoid a progressive biasing of the model by the conversation history. A small percentage of answers might not be correctly formatted (the LLM might respond “The message is conservative” or “Conservative.” or “right-wing” instead of “conservative”), in which case one would simply have to ask the chatbot again. To get an estimate of the accuracy of the labels thus produced, the researcher should ideally have a small subset of the dataset be annotated by experts (typically a few hundreds to a few thousand items of text), so that the labels provided by the LLM can be compared to those provided by experts, which serve as ground truth for the task. The degree of agreement between the two sets of labels can then serve as an estimate of the accuracy of the LLM’s annotations.
The perks of using LLMs to classify text, compared to manual annotation by crowd workers or experts, are immense. It is easy to implement, fast and comparatively inexpensive (thousands of short or medium length texts can be labelled in a few minutes for less than 10$). It is also significantly less time-consuming and easier to set up compared to earlier machine learning-based methods.
@@ 2.1 Limitations of automatic text labelling
As for any method, automatic text annotation using LLMs suffers from some limitations, however, which have not been extensively discussed in the literature. Though early studies are very encouraging, the range of tasks that have been studied so far remains limited; the vast majority of them consists in attributing one of a handful of labels to short- to medium-length texts (typically social media posts or news headlines). Moreover, early results find that accuracy tends to be lower in languages other than English [HCvH24], which is coherent with general observations regarding LLMs’ performances. LLMs are also known to suffer from potential biases, including political ones [LJW+22, MPNR24, ZLS+24]; those could result in systematic errors on certain tasks that would have a greater impact on the research performed on the annotated dataset than the raw accuracy might suggest. As an example, consider an LLM tasked with labelling a set of tweets as Republican-leaning or Democrat-leaning ; if all mislabelled tweets are Republican tweets erroneously labeled as Democrat, the conclusions drawn from the dataset might be wrong, despite the overall accuracy being high.
@@
The issue can also be exacerbated by conscious moderating efforts by the LLM’s creators: various techniques are in place to ensure that the answers provided by LLM-based chat agents avoid using stereotypes seen as offensive [ea23], which can negatively affect their performance on politically sensitive topics like race.
@@
Other problems arise not from the intrinsic nature of LLMs, but from the way in which they are currently being trained and deployed. State-of-the-art LLMs, such as GPT-4, are trained on immense datasets of text scrapped from the web, and regularly updated in a similar fashion. Beyond the possible privacy issues, this has crucial implications regarding the performance of LLMs: they will often perform better on data on which they have been trained than on yet unseen data (e.g., data produced after their training) [Alp10]. As a result, the same LLM might be better at classifying a certain set of tweets than a more recent set it has not yet “seen”. Since the exact nature of the training data is typically not shared with the general public, such phenomena are very hard to predict - see some of our observations in Section 5. In particular, we suspect that some of the most impressive accuracy scores achieved by LLMs on annotation tasks (and reported in the academic literature) might be partially explained by this. This problem, in conjunction with the recurrent updating of models, creates important issues of replicability: results might vary greatly between apparently similar tasks, or when applying different versions of the same model to the same task. Note that this issue can be partially offset by using “frozen” LLMs, i.e., by keeping a copy of a certain version of an LLM at a certain point in time and using it on all tasks.
2.2 Related Works on Prompting for Text Classification
Since the 2000s, sociologists and political scientists have applied machine learning techniques, such as Support Vector Machines, Naive Bayes classifiers, or Maximum Entropy models, to the classification of large volumes of text (e.g. [JD16, NR13, AAS+20, PP11, DMB23, WC17, Alp10]). However, these methods face several limitations [LPL+22], such as struggling with understanding context and informal language patterns common in social media posts, and their performance heavily depends on curated training datasets and domain-specific feature engineering.
A few years ago, Large language models (LLMs) have become available to social scientists and they are greatly facilitating text classification tasks in social sciences [HCvH24, GAK23, Tör23, WR23, LLC+24] (but see [OSMC23] for a more critical perspective). Early results show that off-the-shelves LLMs can perform extremely well, with accuracy levels sometimes equating those of human experts, while being comparatively cheap and easy to deploy. In political discourse analysis in particular, they have been shown to successfully identify political slant, implicit biases, dog whistles, or rhetorical devices that simpler keyword-based approaches would miss [DKG22, HCvH24, GAK23, Tör23, WR23, LLC+24].
The most straightforward way to annotate using LLMs is to ask them the same question, or prompt, as would be asked to human coders [GAK23, AV24, HLG+24, HCvH24]. However, this approach suffers from the models’ sensitivity to the precise wording of the instructions: performances have been observed to improve in response to various “tricks” such as providing the LLM with additional context for the task, giving it examples of correct answers, and asking the LLM to explain its response or to answer while pretending to be a political analyst (see Subsection 4.2). Since then, a literature has emerged documenting such prompt engineering techniques (see e.g., [BMR+20, HZD+21, KGR+22, LDC+22, WWS+22, BG24], and the surveys [SSS+24, VD24]). However, novel prompts tend to be explored in an unsystematic manner and there is no guarantee that a given technique will perform well in an untested setting. This issue can be partially circumvented by annotating part of the data through other means (e.g., human coders) and using this data sample to evaluate the performance of various prompts, before applying the best prompt to the remainder of the data [AV24, WR23]. Doing this is time-consuming, however. It also requires ground truth labels for evaluating prompt performance (which can be costly to obtain), and relies on the researcher’s ability to intuit which candidate formulations would make for better prompts.
As an answer to this problem, a more principled approach has recently been put forward, called Automatic prompt optimization or Automatic prompt engineering, which consists in using LLMs themselves to automatically engineer good prompts [ZMH+22, YWL+24, SDZ23, JPZ+24, CKD24]. We employ and describe such a procedure below, in Subsection 4.3.
3 Prompt Ultra, our automatic dataset labelling app
Though using LLMs to label datasets demands significantly fewer technical skills compared to earlier machine learning-based methods, a basic understanding of coding is still necessary. We are providing a browser app, Prompt Ultra, which can be used to automatically label datasets without any coding being required. It is freely available at https://prompt-ultra.github.io/.
Briefly summarized, one can use the app (and a ChatGPT access key, easily purchasable on their website https://help.openai.com/en/) to label any textual dataset with a prompt of one’s choice. One can also evaluate and compare the performance of various prompts, as well as apply the automatic prompt optimization algorithm from Subsection 4.3. More complete descriptions of the app are available in the Appendix, as well as on the website itself.
@@ 4 Experiments
We evaluate various hand-crafted prompts and automatically optimized prompts (see Subsections 4.2 and 4.3 below) on a range of datasets and classical social sciences tasks using OpenAI’s GPT-3.5 Turbo’s API [YCX+23] (see our code base for details). Our objective is twofold: first, we want to measure the impact of the precise formulation of the prompt on GPT’s accuracy by comparing performance across various prompts. The goal here is not to find some optimal prompt crafting technique that would systematically outperform all others (such panacea is unlikely to exist, as suggested by our experimental results further below); in particular, we do not claim to have examined every possible prompt-crafting trick. Rather, we only want to see whether two reasonable, semantically similar prompts can result in significantly different accuracies on a given task. Second, we want to check whether automatic prompt optimization can help achieve consistently good (though not necessarily optimal) results on all tasks without the need for manual tweaking by the experimenter.
@@ 4.1 Datasets and tasks
A dataset of news articles from major US newspapers [BDSMGN20], with labels representing their political inclination (left, center444Note that the centrist label describes articles that are biased towards a centrist political ideology, and not articles that lack political bias. or right), from which we randomly sample articles.
@@ 4.2 Hand-crafted prompts
As noted in our review of related works, it has been observed that various tricks can help increase LLMs’ performance. As an example, simply adding “Let’s think step by step” at the end of a prompt can result in impressive improvements [KGR+22]. Though there is no systematic theory of “good prompting”, many such tricks have been catalogued (see e.g., [pro]). We have applied several of these standard techniques to craft five different prompts for each task; below are examples from the TweetEval - hate task. The other prompts can be found in the accompanying code.
@@ 4.3 Automatic prompt optimization (APO)
The method is both conceptually simple, and easy to implement; its main a priori downside compared to using hand-crafted prompts is the increased number of calls to the chatbot API required, though the costs remain almost negligible. It can be tested on our free browser-based service, which we further describe in Appendix C.
@@ 5 Results and discussion
We report in Figure 2 the micro-averaged scores on each task of our five types of prompts and of the best prompt generated using automatic prompt generation, as well as their estimated 95% confidence intervals [TYKK22, Gil23]. We observe that the choice of prompt has a significant impact on performance for most tasks: depending on the task, the best prompt results in an observed score that is up to higher than that of the worst prompt. This confirms our claim that careful prompt selection is crucial.
We also note that there is no "miracle" hand-crafted prompt that consistently outperforms all the others. Some heuristics and specific wordings work well on certain tasks, and poorly on others. This is in line with observations made in the literature; as an example, adding “Let’s think step by step” at the end of a prompt is shown to outperform adding the semantically similar “Let’s work this out in a step by step way to be sure we have the right answer” for the task studied in [YWL+24] (see Table 1 thereof), while the converse is true in [ZMH+22] (Table 7).
@@
We now turn to the specific performances of Automatic prompt optimization (APO). APO beats hand-crafted prompts at four tasks; among the remaining three, it is second best for TE-sent (with a score nearly equal to that of the best hand-crafted prompt), and is roughly at the level of the median of the hand-crafted prompts for TE-off and TE-hate. This shows that automatically optimizing prompts results in consistently good performance without the need for prior testing of various hand-crafted prompts on a subset of the dataset of interest. Note that the expected performance of APO could be improved by increasing the number of samples on which each generation is tested (the samples mentioned in Subsection 4.3), in exchange for an increase in computational costs.
Interestingly, and in line with the remarks of the previous paragraph, the final prompts generated using APO do not differ much from the ones used to start the optimization process, despite the often large gap between their associated accuracies555We report the best prompt returned by the optimization process for each task in the Appendix.. For TE-hate, for example, the initial prompt is “Classify the following message as hateful if it contains hate speech or non-hateful if it does not contain hate speech. Output only “hateful” or “non-hateful” without quotes” and the automatically optimized prompt is “Check for hate speech in the following message to determine if it is hateful, then classify it as either "hateful" or "non-hateful"”.
@@
We have observed a curious phenomenon: as shown in Figure 3, all prompts result in much higher score when tested on the training set of the TE-hate dataset than on its test set (where “training” and “test” refer to the split made by the creators of the dataset [BBF+19]). This is a priori surprising, as both subsets should be drawn from the same distribution, and as such our prompts should not perform any better on the training set than on the test set666Note that there is no issue of overfitting here, as the prompts were not “trained” in any way on what we call the training set in this context.. Our explanation is that the training set of TE-hate is part of the enormous amount of data on which GPT-3.5 Turbo, the model which we used, has been trained; as a result, it performs better on it than on the test set (which was either not included, or included in a different fashion). As the exact data on which the model was trained was not made public, we cannot confirm this hypothesis with absolute certainty. As a result, we can say that ChatGPT is “cheating” when labelling the training set of TE-hate, in the sense that its performance is superior to what it would be on a similar dataset to which it did not have access during its training (e.g. data that was produced after its training). This explains why we used the test splits of our datasets whenever possible, in the hope that as (we assume it to be the case) for TE-hate, they have not been included in the training set of the model, making for a fairer assessment of the capabilities of the method.
@@
This problem makes it hard to predict future performances based on past experiments, as we have no guarantee as to what was and was not used in the models’ training. Note for example that a similar gap in performance between the training set and the test set is not observable for TE-emotion. Ideally, one should only test the models on data that was produced after their training. This could however prove difficult and overly restrictive: curating useful datasets is a challenging and time-consuming task in itself, and most datasets available online are at least a few years old, thereby predating all last generation LLMs.
@@
As a side note, this phenomenon could explain why we have observed slightly weaker performances on average than those reported in recent literature [HCvH24, GAK23, Tör23] on similar tasks: we suspect that the authors of these articles might not have been aware of this pitfall, and failed to take measures to circumvent it. It could also be due to intrinsic differences in difficulty between the datasets used, as we tried to include both easy and hard datasets for diversity (in particular, it has been noted before, e.g., in [HCvH24], that performance drops for very long texts, such as the articles from the AS-pol dataset). We cannot verify this hypothesis, as (most of) the datasets used in the references that we cite are not available anymore due to a policy change from the social media platform X.
A note on our choice of datasets and task complexity
We selected on purpose tasks that are commonly performed by social scientists and natural language processing researchers so as to make the relevance of our claims (variation in prompt performance and the benefits of automatic optimization) immediately apparent to both communities. A downside of this choice is that though distinct, the tasks considered are somewhat similar and inevitably simple. Another limitation of our datasets is that though some tasks are harder than others, none are hard enough to present a real challenge to a well-informed human being. In future work, it would be interesting to measure the accuracy and sensitivity of LLMs to the prompts chosen both on a greater variety of tasks and in limit cases where even human experts might struggle. As a final remark, note that due to the very nature of the tasks at hand, some labels in our datasets are not unambiguously correct: it is for instance easy to find tweets in the TE-emotion dataset that could be classified as either joyful or optimistic. This fact, while not unique to our datasets, necessarily has an impact on the performances reported.